OpenStudy (anonymous):

If 2/3p + 6 = 7/6p, what is the value of p?

OpenStudy (anonymous):

Any tips will he helpful thank you

OpenStudy (anonymous):

\[\frac{2}{3}p+6=\frac{7}{6}p\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok that was wrong, sorry multiply both sides by 6 and start with \[4p+36=7p\] now it takes only two steps

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what should I do next?

OpenStudy (anonymous):

Now you just need to combine like terms.

OpenStudy (anonymous):

Bring the 4p over to the 7p by subtracting both sides by 4p.

OpenStudy (anonymous):

so is it 3p?

OpenStudy (anonymous):

Yes on the right side it's 3p. Now the equation looks like this:\[36=3p\]

OpenStudy (anonymous):

Do you know what to do then?

OpenStudy (anonymous):

are you supposed to fing the p?

OpenStudy (anonymous):

Yes lol that was what we were doing in the first place. :)

OpenStudy (anonymous):

Divide both sides by 3 so you can get p by itself.

OpenStudy (anonymous):

so divide 3p by 3?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

is that 1

OpenStudy (anonymous):

Well it would equal 1p since the p was never cancelled out when you divided by 3. Now divide 36 by 3.

OpenStudy (anonymous):

that's 12

OpenStudy (anonymous):

Yes that is right! Now you have:\[12=p\]

OpenStudy (anonymous):

There's how you get it! :)

OpenStudy (anonymous):

oh ok,Thanks1

OpenStudy (anonymous):

how to I GIVE OUT MEADELS?

OpenStudy (anonymous):

Just click on the best response button on the side of my comment. And you're welcome!